OpenStudy (anonymous):

You accidentally throw your car keys horizontally at 8.0 m/s from a cliff 64-m high. How far from the base of the cliff should you look for the keys?

OpenStudy (curiouser_alice):

How long is it in the air (use gravity equation) x 8 m/s (horizontal speed). Force of gravity on earth = 9.8 m/s2.

OpenStudy (curiouser_alice):

Horizontal motion and vertical motion are independent

OpenStudy (anonymous):

29 m (3.61 s to fall)

OpenStudy (anonymous):

Its 28.56 m Here is the explation, As \[ y=u \sin \theta-\frac{ 1 }{ 2 } g t{2} \] where y=-64 (for downward motion), u=8m/s and for horizontal throw, \[\theta=0^{0}\] hence t=3.57s again applying, \[x=u \cos \theta * t\] we have x=28.56m

OpenStudy (anonymous):

actually, it us 3.55 seconds

OpenStudy (anonymous):

so the answer is 28.4 meters